# Highest–weight truncation, graded EFT structure, and renormalization of black hole Love numbers

### v2

@@ Abstract
The static tidal Love numbers of four–dimensional black holes vanish identically, unlike their nontrivial dynamical response at finite frequency. Recent work has provided three complementary descriptions of this phenomenon: an emergent $\mathrm{SL}(2,\mathbb{R})$ organization of static near–zone perturbations, a graded logarithmic and multi–zeta structure in Shell Effective Field Theory (Shell EFT), and an on–shell matching framework based on gravitational Raman scattering with renormalization group (RG) running. We show that these features arise from a common near–zone truncation mechanism. For a massless scalar field, horizon regularity selects a unique static solution forming a highest–weight–type representation, truncating the hypergeometric solution to a finite polynomial and eliminating the independent decaying branch at large radius. This excludes a static Wilson coefficient in the effective theory. We demonstrate that the same truncation operates in the static Regge–Wheeler and Zerilli equations for four–dimensional Schwarzschild black holes. Analytic continuation of the horizon–regular solution to small frequency—via the Coulomb–hypergeometric or Mano–Suzuki–Takasugi formalisms—preserves this truncation as an anchoring condition for the renormalized angular momentum parameter. The resulting low–frequency expansion is controlled by Gamma and hypergeometric functions, generating a graded algebra of logarithms and odd Riemann zeta values. Within this structure no invariant of negative weight exists in the static sector, so the vanishing of the static Love number follows as a structural consequence. This explains the “zero–sum” rule of Shell EFT and why the self–induced RG flow in gravitational Raman scattering cannot generate a static invariant.
@@ 1 Introduction
In four–dimensional general relativity, black holes behave in a strikingly different manner. For asymptotically flat black holes, the static tidal Love numbers vanish identically, both for Schwarzschild and Kerr spacetimes Landry and Poisson (2015); Chia (2021); Le Tiec et al. (2021); Le Tiec and Casals (2021); Hui et al. (2022); Ben Achour et al. (2022); Charalambous et al. (2021); Perry and Rodriguez (2023); Riva et al. (2024); Combaluzier-Szteinsznaider et al. (2025); Gounis et al. (2025); Kehagias and Riotto (2025). This result has been established through direct perturbative calculations and holds independently of spin. Moreover, the robustness of the vanishing has been tested beyond linear response. For scalar perturbations with self-interactions, it was shown that while generic nonlinear interactions can generate tails, the tidal Love numbers vanish to all orders for nonlinear sigma models, highlighting a structural protection mechanism analogous to that in General Relativity De Luca et al. (2023). This strengthens the conclusion that the absence of static response is not an artifact of linearization but a structural property of four-dimensional black holes in general relativity.
At the same time, black holes do exhibit a nontrivial dynamical tidal response: at finite frequency the response coefficients are nonzero Perry and Rodriguez (2023); Saketh et al. (2024); Combaluzier–Szteinsznaider et al. (2025); Kosmopoulos et al. (2025); Charalambous et al. (2021); Chia (2021); Perry and Rodriguez (2025); Chakraborty et al. (2025) and display a rich frequency dependence. In particular, the dynamical Love numbers exhibit a characteristic running behavior whose coefficient can be computed explicitly and contributes to the binary black-hole gravitational waveform at high post-Newtonian order (8PN) Chakraborty et al. (2025). The coexistence of vanishing static response with nontrivial dynamical behavior poses a conceptual puzzle.
@@
Third, gravitational Raman scattering has established, at the level of on–shell amplitudes and renormalization group flow, that matching to Schwarzschild phase shifts requires the absence of a static Wilson coefficient, with the dynamical response exhibiting a distinctive self–induced running structure Ivanov et al. (2024).
Previous works identified these structures independently. What has not been shown is that the graded multi–zeta algebra of Shell EFT and the Raman RG structure arise as rigid consequences of highest–weight truncation in the near zone. The purpose of this work is to establish this structural connection. We demonstrate explicitly, using a massless scalar field as a controlled setting, that the highest–weight condition enforced by horizon regularity acts as a selection rule excluding an independent static response parameter. When translated into the EFT language, this selection rule directly implies the absence of the corresponding static Wilson coefficient and explains why matching to Schwarzschild phase shifts in Raman scattering requires $\lambda_{\ell,0}=0$. We then show that, once the exterior solution is expressed in the Coulomb–hypergeometric basis employed in Shell EFT, analytic continuation of the highest–weight static solution to finite frequency forces the remaining response coefficients to organize into a graded algebra of logarithms and zeta values. In this sense, neither the Shell EFT graded structure nor the Raman renormalization group constraints are independent or accidental features of matching. Rather, they are the effective–theory imprints of the emergent near–zone highest–weight symmetry. By making this connection explicit, we provide a unified explanation for the vanishing of static black–hole Love numbers and the highly constrained structure of their dynamical tidal response.
@@ 7 Connection to Gravitational Raman Scattering and Renormalization Group Structure
In this work we have clarified the conceptual origin of the vanishing static Love numbers of four–dimensional black holes by explicitly connecting three apparently distinct perspectives: the structure of the static near–zone perturbation equations, the analytic organization of tidal response coefficients revealed by Shell Effective Field Theory (Shell EFT), and the on–shell matching framework based on gravitational Raman scattering.
@@
Taken together, these results demonstrate that the Shell EFT structure and the Raman scattering RG flow are not independent or accidental features of the matching procedure. Rather, they can be understood as the effective–theory imprint of horizon regularity and the associated highest–weight structure of the static black–hole perturbation problem. The uniqueness of the static solution eliminates the static invariant, while the Coulomb–hypergeometric realization controls the detailed transcendental content and scale dependence of the allowed dynamical response. Symmetry excludes forbidden operators, analyticity organizes the allowed coefficients, and on–shell matching encodes their renormalization.
@@
Although our detailed analysis focused on a massless scalar field for clarity, the underlying logic is more general. We have argued that the same interplay between near–zone structure, horizon regularity, analytic continuation, and on–shell matching governs gravitational perturbations of four–dimensional Schwarzschild black holes, and is consistent with the known vanishing of static gravitational Love numbers. Extending this framework to rotating backgrounds, where additional physical effects such as superradiance and mode coupling arise, represents a natural and important direction for future work.

### v3

@@ Abstract
The static tidal Love numbers of four–dimensional black holes vanish identically, unlike their nontrivial dynamical response at finite frequency. Recent work has provided three complementary descriptions of this phenomenon: an emergent $\mathrm{SL}(2,\mathbb{R})$ organization of static near–zone perturbations, a graded logarithmic and multi–zeta structure in Shell Effective Field Theory (Shell EFT), and an on–shell matching framework based on gravitational Raman scattering with renormalization group (RG) running. We show that these features arise from a common near–zone truncation mechanism. For a massless scalar field, horizon regularity selects a unique static solution forming a highest–weight–type representation, truncating the hypergeometric solution to a finite polynomial and eliminating the independent decaying branch at large radius. This excludes a static Wilson coefficient in the effective theory. We show that the same truncation operates in the static Regge–Wheeler and Zerilli equations for four–dimensional Schwarzschild black holes. Analytic continuation of the horizon–regular solution to small frequency—via the Coulomb–hypergeometric or Mano–Suzuki–Takasugi formalisms—preserves this truncation as an anchoring condition for the renormalized angular momentum parameter. The resulting low–frequency expansion is controlled by Gamma and hypergeometric functions, generating a graded algebra of logarithms and odd Riemann zeta values. Within this structure, no invariant of negative weight exists in the static sector, so the vanishing of the static Love number follows as a structural consequence. This explains the “zero–sum” rule of Shell EFT and why the self–induced RG flow in gravitational Raman scattering does not generate a static invariant.
@@ 1 Introduction
In four–dimensional general relativity, black holes behave in a strikingly different manner. For asymptotically flat black holes, the static tidal Love numbers vanish identically, both for Schwarzschild and Kerr spacetimes Landry and Poisson (2015); Chia (2021); Le Tiec et al. (2021); Le Tiec and Casals (2021); Hui et al. (2022); Ben Achour et al. (2022); Charalambous et al. (2021); Perry and Rodriguez (2023); Riva et al. (2024); Combaluzier-Szteinsznaider et al. (2025); Gounis et al. (2025); Kehagias and Riotto (2025); Bhatt et al. (2023). This result has been established through direct perturbative calculations and holds independently of spin. Moreover, the robustness of the vanishing has been tested beyond linear response. For scalar perturbations with self-interactions, it was shown that while generic nonlinear interactions can generate tails, the tidal Love numbers vanish to all orders for nonlinear sigma models, highlighting a structural protection mechanism analogous to that in General Relativity De Luca et al. (2023). This strengthens the conclusion that the absence of static response is not an artifact of linearization but a structural property of four-dimensional black holes in general relativity.
At the same time, black holes do exhibit a nontrivial dynamical tidal response: at finite frequency the response coefficients are nonzero Perry and Rodriguez (2023); Saketh et al. (2024); Combaluzier–Szteinsznaider et al. (2025); Kosmopoulos et al. (2025); Charalambous et al. (2021); Chia (2021); Perry and Rodriguez (2025); Chakraborty et al. (2025); Bhatt et al. (2025) and display a rich frequency dependence. In particular, the dynamical Love numbers exhibit a characteristic running behavior whose coefficient can be computed explicitly and contributes to the binary black-hole gravitational waveform at high post-Newtonian order (8PN) Chakraborty et al. (2025). The coexistence of vanishing static response with nontrivial dynamical behavior poses a conceptual puzzle.
@@
Third, gravitational Raman scattering has been developed in a series of works Ivanov et al. (2024); Correia and Isabella (2025); Caron-Huot et al. (2025); Ivanov et al. (2026), establishing, at the level of on–shell amplitudes and renormalization group flow, that matching to Schwarzschild phase shifts requires the absence of a static Wilson coefficient, with the dynamical response exhibiting a distinctive self–induced running structure.
Previous works identified these structures independently. What has not been shown is that the graded multi–zeta algebra of Shell EFT and the Raman RG structure can arise from highest–weight truncation in the near zone. The purpose of this work is to establish this structural connection. We demonstrate explicitly, using a massless scalar field as a controlled setting, that the highest–weight condition enforced by horizon regularity acts as a selection rule excluding an independent static response parameter. When translated into the EFT language, this selection rule directly implies the absence of the corresponding static Wilson coefficient and explains why matching to Schwarzschild phase shifts in Raman scattering requires $\lambda_{\ell,0}=0$. We then show that, once the exterior solution is expressed in the Coulomb–hypergeometric basis employed in Shell EFT, analytic continuation of the highest–weight static solution to finite frequency forces the remaining response coefficients to organize into a graded algebra of logarithms and zeta values. In this sense, neither the Shell EFT graded structure nor the Raman renormalization group constraints are independent or accidental features of matching. Rather, they are the effective–theory imprints of the emergent near–zone highest–weight symmetry. By making this connection explicit, we provide a unified explanation on the vanishing of static black–hole Love numbers and the highly constrained structure of their dynamical tidal response.
@@ 7 Connection to Gravitational Raman Scattering and Renormalization Group Structure
In this work we have clarified a coherent perspective on the conceptual origin of the vanishing static Love numbers of four–dimensional black holes by explicitly connecting three apparently distinct perspectives: the structure of the static near–zone perturbation equations, the analytic organization of tidal response coefficients revealed by Shell Effective Field Theory (Shell EFT), and the on–shell matching framework based on gravitational Raman scattering.
@@
Taken together, these results demonstrate that the Shell EFT structure and the Raman scattering RG flow are not independent or accidental features of the matching procedure. Rather, they can be understood as the effective–theory imprint of horizon regularity and the associated highest–weight structure of the static black–hole perturbation problem. The uniqueness of the static solution eliminates the static invariant, while the Coulomb–hypergeometric realization controls the detailed transcendental content and scale dependence of the allowed dynamical response. In this sense, the Schwarzschild case suggests an interpretation in which symmetry excludes forbidden operators, analyticity organizes the allowed coefficients, and on–shell matching encodes their renormalization.
@@
Although our detailed analysis focused on a massless scalar field for clarity, the underlying logic is more general. We have argued that the same interplay between near–zone structure, horizon regularity, analytic continuation, and on–shell matching governs gravitational perturbations of four–dimensional Schwarzschild black holes, and is consistent with the known vanishing of static gravitational Love numbers. However, it remains an open question to what extent this structure extends to more general spacetimes, such as rotating black holes, where additional effects may arise. Extending this framework to rotating backgrounds, where additional physical effects such as superradiance and mode coupling arise, represents a natural and important direction for future work.